OpenStudy (anonymous):

Solve the system of equations using the addition method. 4x-5y=1 3x=6y-4

OpenStudy (anonymous):

@Agent47

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (agent47):

first put the unknowns on the same side: 4x-5y=1 3x-6y=-4

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

first get them to look like each other, by this I mean x+y = something so 4x-5y = 1.....I jsut noticed agent doing the same thing.... listen to him.

OpenStudy (agent47):

zzr0ck can finish this :P

OpenStudy (anonymous):

No Agent I like u will u finish it

OpenStudy (agent47):

lol KE i don't like u.. jk jk ok i will

OpenStudy (anonymous):

lol :)

OpenStudy (agent47):

So add the two equations: 4x-5y=1 3x-6y=-4 --------- 7x-11y=-3

OpenStudy (agent47):

but i dont see how using the addition method helps here...

OpenStudy (anonymous):

It says to simplify the answer and write as a ordered pair

OpenStudy (agent47):

wait disregard all the previous steps..

OpenStudy (agent47):

starting over: first put the unknowns on the same side: 4x-5y=1 3x-6y=-4 multiply the second equation by -1 throughout: 4x-5y=1 -3x+6y=4

OpenStudy (agent47):

now add the two equations: x+y=3

OpenStudy (agent47):

x=3-y 4x-5y=1; 12-4y-5y=1; 12-9y=1; -9y=-11; y=11/9... am i doing something totally wrong? zzr0ck can u go over my work?

OpenStudy (agent47):

x=3 - 11/9 = 16/9

OpenStudy (anonymous):

there is also a option for infintenty amount of solutions or no solution

OpenStudy (agent47):

as an ordered pair: (16/9;11/9) can u post all choices?

OpenStudy (anonymous):

A: The solution is ___ B: There are infintety many solutions c: there is no solution

OpenStudy (agent47):

ok im making a mistake here, i just tested the answer, it's wrong..

OpenStudy (anonymous):

I'm lost also @Agent47

OpenStudy (agent47):

(26/9 ;19/9) can u check this?

OpenStudy (anonymous):

ah ill try from your previous equations, 4x - 5y = 1 Equation 1 3x - 6y = -4 Equation 2 Ill multiply equation 1 by 3 and multiply equation 2 by 4 to get 12x - 15y = 3 12x - 24y = -16 subtracting from each other we get 9y = 19 y = 19/9 substituting this value into Equation 1 4x - 5*(19/9) = 1 4x = (9/9) + (95/9) = 104/9 x = 26/9 i think.. let me recheck

OpenStudy (agent47):

oh there, i think this time we were both right with shi

OpenStudy (anonymous):

ok does it matter which one goes first the 19/9 or 26/9?

OpenStudy (agent47):

26/9 goes first

OpenStudy (agent47):

ordered pair: x; y

OpenStudy (anonymous):

yeah i plugged them in, they are correct ordered pairs are usually written (x,y) or you can just say x = this y = that

OpenStudy (anonymous):

Thank u sooo much that was correct